OpenStudy (anikate):

find the center and radius of (x+10)^2 +(y-1)^2 = 9

OpenStudy (anonymous):

Center: (-10, 1) Radius: 3

OpenStudy (anonymous):

Do you know why?

OpenStudy (anikate):

nope plz explain

OpenStudy (anikate):

brb afk

OpenStudy (anonymous):

Formula for a circle: \[(x-h)^{2}+(y-k)^{2}=r ^{2}\] (h, k)=center r=radius

OpenStudy (anikate):

wouldn't it be (-10, -1)= center

OpenStudy (anonymous):

No because in the formula it is (y-k), in the question it is also (y-1) so the 1 stays positive the 10 is negative because the formula is (x-h) and the question shows (x+10). Do you understand?

OpenStudy (anonymous):

I know it's kind of confusing the way I explained it...

OpenStudy (anikate):

Oh ok i c thx J.L

OpenStudy (anonymous):

No problem :) let me know if you need anymore help

OpenStudy (asnaseer):

it can be made clearer if you rewrite the equation in standard form as follows:\[(x-(-10))^2 +(y-(1))^2 = 3^2\]

OpenStudy (anikate):

OK THX :D

OpenStudy (anikate):

thx too asnaseer

OpenStudy (asnaseer):

yw :)